OpenStudy (anonymous):

If f(x)=x+1 and g(X)=x^2-1, for what values of x does g[f(x)]=0?

OpenStudy (anonymous):

(x+1)^2 - 1 =0 (x+1)^2= 1 so x+1=+/- 1 so x=0 orx=-2

OpenStudy (anonymous):

\[g(f(x))=g(x+1)=(x+1)^2-1=x^2+2x+1-1=x^2+2x=x(x+2)\] so it is zero if \[x=0,x=-2\] better answer is from shnarayana